write an equation in point slope form that passes through (8,-2)and is parellel to the graph of y= 2x-4.
Answer by Earlsdon(6098)

Recall that the point=slope form is:
y = mx + b
You can find the slope of the new line when you remember that parallel lines have the same slope. The slope of the given line is easily seen when you compare the given equation with the general slope-intercept form.
y = 2x - 4 compare with:
y = mx + b and you can see that the slope of the given line, and of any line parallel to it, is m = 2
So you can start your new equation off with:
y = 2x + b Now, since this line passes through the point (8, -2), then the coordintates of the point (x = 8 and y = -2) must satisfy this equation. Substitute these values of x and y into you new equation:
-2 = 2(8) + b now you can solve this for b.
-2 = 16 + b Subtract 16 from both sides of the equation.
-18 = b Now you can write your final equation:
y = 2x - 18